OpenStudy (anonymous):

A construction company needs 2 weeks to construct a family room and 3 days to add a porch. Find the ratio of the time it takes for constructing the porch to the time constructing the family room, with all units in weeks. (Show steps, please and thank you.)

OpenStudy (vishal_kothari):

TIME TAKEN FOR PORCH/TIME TAKEN FOR ROOM=3 WEEKS/2WEEKS

OpenStudy (anonymous):

Porch: 3 days = 3/7 weeks Room: 2 weeks So, ratio: 3/7 / 2 = 3/14

OpenStudy (anonymous):

Thank you everyone :)